# Unavoidable Canonical Nonlinearity Induced by Gaussian Measures Discretization

### v3

@@ Abstract
When we consider canonical averages for classical discrete systems, typically referred to as substitutional alloys, the map $\phi$ from many-body interatomic interactions to thermodynamic equilibrium configurations generally exhibits complicated nonlinearity. This canonical nonlinearity is fundamentally rooted in deviations of the discrete configurational density of states (CDOS) from continuous Gaussian families, and has conventionally been characterized by the Kullback-Leibler (KL) divergence on discrete statistical manifold. Thus, the previous works inevitablly missed intrinsic nonlinearities induced by discretization of Gaussian families, which remains invisible within conventional information-geometric descriptions. In the present work, we identify and quantify such unavoidable canonical nonlinearity by employing the 2-Wasserstein distance with a cost function aligned with the Fisher metric for Gaussian families. We derive an explicit expression for the Wasserstein distance in the limit of vanishing discretization scale $d\to 0$: $W_{2}=d\sqrt{\textrm{Tr}\left(\Gamma^{-1}\right)/12}$, where $\Gamma$ denotes covariance matrix of the CDOS. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems. In addition, we confirm that this $W_{2}$-KL equivalence admits a natural generalization beyond Gaussian families, provided that the transport cost is aligned with the Fisher metric of an underlying statistical submanifold and the discretization scale induces infinitesimal parameter variations.
@@ I Introduction
Existing approaches to quantify such canonical nonlinearity are primarily based on information geometry, in particular the Kullback-Leibler (KL) divergence:klo The nonlinearity was evaluated by comparing the discrete CDOS of a real system with a reference Gaussian distribution having the same mean and covariance matrix, where the Gaussian is discretized on the same configurational support as the real CDOS. While this framework captures deviations of the CDOS from Gaussian, it necessarily includes additional contributions arising from the discretization of continuous Gaussian families themselves. Such discretization-induced effects remain invisible on discrete statistical manifolds, where probability measures are compared only on a fixed set of discrete supports.
In the present work, we identify this contribution as an unavoidable canonical nonlinearity (UCN). To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization. Motivated by this requirement, we employ optimal transport theory, specifically the 2-Wasserstein distance $W_{2}$. By introducing a quadratic cost function aligned with the Fisher metric of Gaussian families, the Wasserstein distance becomes directly compatible with information-geometric descriptions while remaining sensitive to discretization-induced geometric distortions. We derive an explicit expression for $W_{2}$ between a continuous Gaussian and its discretized counterpart in the limit of vanishing discretization scale $d\to 0$, yielding a universal form depending only on the covariance matrix of the CDOS and $d$. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone.
@@ II Concept and Derivation
Based on this observation, previous studies have quantified canonical nonlinearity by measuring the difference between $P$ and $P_{c}$ using the KL divergence. Since the KL divergence requires the two distributions to be defined on the same support (or to satisfy an inclusion relationship of supports), the continuous Gaussian $P_{c}$ has been discretized on the same configurational support as $P$, yielding a discretized Gaussian distribution $P_{d}$. Accordingly, canonical nonlinearity has been characterized by the KL divergence $D\left(P:P_{d}\right)$, which extends to the KL divergence between the canonical distributions induced by the CDOSs $P$ and $P_{d}$.
@@
Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$. Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$. Explicitly, this process is expressed as
@@
$\lim_{d\to 0}W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right)$
$=$ $2\cdot\mathbb{E}_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right].$
Eq.(II) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure. This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold. These contributions are invisible to existing KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
In a realistic discrete system on a lattice, the following three features are unavoidable: (i) $d$ takes a nonzero positive value, (ii) the domain of the CDOS is bounded, and (iii) the domain is generally asymmetric. All of these features are dominated by the underlying lattice. Therefore, to address the effect of UCN in realistic systems, numerical approaches, such as systematic comparisons between Eq. (II) and Eq. (8) under various conditions, are essential for disentangling the individual contributions of these three lattice-induced effects.
@@
where $f$ corresponds to the number of effective degrees of freedom relevant to the description of the system interested (and, subsequently, to the transport problem). Then we consider the problem to measure the difference between function $P\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families. We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant. More generally, our statement rely on the locally-constant character of $\Omega$ under infinitesimal variations in $\xi$. Then we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P:P^{\prime}\right)$ as
@@
Then, we discretize the underlying continuous space for $P$, by using $f$-dimensional hypercubic with side length $d$, and consider its limit of $d\to 0$. At the same time, we assume that the uncertainty or variation of the parameters $\xi$, induced by the discretization scales $d$, can be given by
@@
where $a>0$ denotes a scalar constant, typically characterizing the sensitivity of the parameters $\xi$ to the discretization scale. Then we interpreted $\delta\xi$ as the following i.i.d. random vector with its distribution taking $\rho$:
@@
$W_{2}^{2}\left(P:P^{\prime}\right)=\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
@@
The proportionality constant in Eq. (25) explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation. For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$ and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed. In more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement: the discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions. We emphasize that the above leading-order correspondence relies on the assumption that the discretization-induced transport cost is dominated by local fluctuations: This condition can be naturally satisfied for distributions with sufficiently fast-decaying tails (including Gaussian families), where distant contributions to the quadratic transport cost are reasonably suppressed. In contrast, for distributions exhibiting heavy tails or slow decays, non-local contributions from the tail region may modify the scaling behavior of the $W_{2}$ distance in the limit $d\to 0$, potentially requiring corrections including the higher-order terms in $d$. The present analysis therefore characterizes a universal local correspondence, whose extension to heavy-tailed distributions is left for our future work.
@@
Through the discussions about generalization, we now see that the extension of $\delta\xi$ to random vector is not merely a matter of average, but is structurally necessary: The trace structure appearing in the $W_{2}$ of Eq. (24) can only emerge in quadratic form of the KL divergence, if the outer product of $\delta\xi$ is effectively proportional to the identity matrix $I$, i.e., $\delta\xi\delta\xi^{\mathsf{T}}\propto d^{2}I_{f}$. Since the outer product of a single vector is necessarily rank one, while the required structure is full rank with $f$ degrees of freedom, an extension to a probabilistic ensemble of $\delta\xi$ is unavoidable for systems with $f\geq 2$.
@@
Finally, we comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence. While the KL divergence itself is invariant under reparametrizations on the statistical manifold, Eqs. (24)-(25) suggest that its expectation under $\rho$ depends on the trace of Fisher metric $\Omega$. This does not represent a contradiction, but can be reasonably understood from the following two aspects. (i) Taking the expectation $\mathbb{E}_{\rho}$ in Eq. (25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector (the associated probability density $\rho$ typically reflects a modeling choice related to discretization or coarse-graining, depending on the problem under consideration). (ii) In general, when higher-order terms are present, the KL divergence expanded up to second order in $\delta\xi$, yielding a quadratic form, is explicitly governed by the Fisher metric $\Omega$ associated with the chosen parameter $\xi$. In such cases, once combined with the discretization scale and the extrinsic expectation over parameter variations, the resulting numerical value depends on the chosen parameter and Fisher metric. From these viewpoints, the explicit expression $d^{2}\mathrm{Tr}(\Omega)/12$ generally depends on the chosen parameter, yet the correspondence between the 2-Wasserstein transport cost and the expected KL divergence should be understood as an invariant structural relation. In particular, for the case of UCN in this work, since the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and the parameter variations scale linearly with the discretization length, the proportionality constant can be fixed. As a result, the correspondence yields an equality, Eq. (II), rather than a mere proportionality. In more general settings, different parameters and modeling choices for parameter variations lead to different numerical prefactors. Nevertheless, the universal correspondence $W_{2}^{2}\propto 2/a^{2}\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs.

### v4

@@ Abstract
When we consider canonical averages for classical discrete systems, typically referred to as substitutional alloys, the map $\phi$ from many-body interatomic interactions to thermodynamic equilibrium configurations generally exhibits complicated nonlinearity. This canonical nonlinearity is fundamentally rooted in deviations of the discrete configurational density of states (CDOS) from continuous Gaussian families, and has conventionally been characterized by the Kullback-Leibler (KL) divergence on discrete statistical manifold. Thus, the previous works inevitablly missed intrinsic nonlinearities induced by discretization of Gaussian families, which remains invisible within conventional information-geometric descriptions. In the present work, we identify and quantify such unavoidable canonical nonlinearity by employing the 2-Wasserstein distance with a cost function aligned with the Fisher metric for Gaussian families. We derive an explicit expression for the Wasserstein distance in the limit of vanishing discretization scale $d\to 0$: $W_{2}=d\sqrt{\textrm{Tr}\left(\Gamma^{-1}\right)/12}$, where $\Gamma$ denotes covariance matrix of the CDOS. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems. In addition, we confirm that this $W_{2}$-KL equivalence admits a natural generalization beyond Gaussian families, provided that the transport cost is aligned with the Fisher metric of an underlying statistical submanifold and the discretization scale links to infinitesimal parameter variations.
@@ I Introduction
Existing approaches to quantify such canonical nonlinearity are primarily based on information geometry, in particular the Kullback-Leibler (KL) divergence:klo The nonlinearity was evaluated by comparing the discrete CDOS of a real system with a reference Gaussian distribution having the same mean and covariance matrix, where the Gaussian is discretized on the same configurational support as the real CDOS. While this framework captures deviations of the CDOS from Gaussian, it necessarily includes additional contributions arising from the discretization of continuous Gaussian families themselves. Such discretization-induced effects remain invisible on discrete statistical manifolds, where probability measures are compared only on a fixed set of discrete supports. This limitation ultimately originates from the absence of a framework that allows one to interpret discretization-induced geometric distortions of continuous distributions within information geometry.
In the present work, we define the discretization-induced geometric contribution for Gausian families as an unavoidable canonical nonlinearity (UCN), and demonstrate that it admits a consistent information-geometric interpretation, thereby addressing this missing piece. To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization. Motivated by this requirement, we employ optimal transport theory, specifically the 2-Wasserstein distance $W_{2}$. By introducing a quadratic cost function aligned with the Fisher metric of Gaussian families, the Wasserstein distance becomes directly compatible with information-geometric descriptions while remaining sensitive to discretization-induced geometric distortions. We derive an explicit expression for $W_{2}$ between a continuous Gaussian and its discretized counterpart in the limit of vanishing discretization scale $d\to 0$, yielding a universal form depending only on the covariance matrix of the CDOS and $d$. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone.
@@ II Concept and Derivation
Based on this observation, previous studies have quantified canonical nonlinearity by measuring the difference between $P$ and $P_{c}$ using the KL divergence. Since the KL divergence requires the two distributions to be defined on the same support (or to satisfy an inclusion relationship of supports), the continuous Gaussian $P_{c}$ has been discretized on the same configurational support as $P$, yielding a discretized Gaussian distribution $P_{d}$. Accordingly, canonical nonlinearity has been characterized by the KL divergence $D\left(P:P_{d}\right)$, which subsequently extends to the KL divergence between the canonical distributions induced by the CDOSs $P$ and $P_{d}$.
@@
Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with mean vector $\mu$ and covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$. Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$. Explicitly, this process is expressed as
@@
$\lim_{d\to 0}W_{2}^{2}\left(P_{c}\left(\mu,\Gamma\right):P_{d}\right)$ $=$ $2\cdot\mathbb{E}_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right]$ (17)
$=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).\quad$
Eq.(17) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure. This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold. These contributions are invisible to existing KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
In a realistic discrete system on a lattice, the following three features are unavoidable: (i) $d$ takes a nonzero positive value, (ii) the domain of the CDOS is bounded, and (iii) the domain is generally asymmetric. All of these features are dominated by the underlying lattice. Therefore, to address the effect of UCN in realistic systems, numerical approaches, such as systematic comparisons between Eq. (17) and Eq. (8) under various conditions, are essential for disentangling the individual contributions of these three lattice-induced effects.
@@
where $f$ corresponds to the number of effective degrees of freedom relevant to the description of the system interested (and, subsequently, to the transport problem). Then we consider the problem to measure the difference between function $P_{\xi}\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families. We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant. More generally, our statement rely on the locally-constant character of $\Omega$ under infinitesimal variations in $\xi$. Now we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P_{\xi}:P^{\prime}\right)$ as
@@
Then, we discretize the underlying continuous space for $P$, by using $f$-dimensional hypercubic with side length $d$, and consider its limit of $d\to 0$. Under this discretization, we introduce the following linear relationship betwen $\delta\xi$ and $d$:
@@
where $a>0$ reflects the relative scaling arising from the identification of the local quadratic structure of the KL divergence with the discretization-induced transport cost. Then we extend $\delta\xi$ to the following i.i.d. random vector with its distribution taking $\rho$:
@@
$W_{2}^{2}\left(P_{\xi}:P^{\prime}\right)=\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
@@
While the W2-KL correspondence in Eq. (17) provides a transparent physical intuition, i.e., the transport cost between continuous and discrete Gaussian is interpreted as the expected parallel translation of continuous Gaussian, in more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement. The discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions. We emphasize that the above leading-order correspondence relies on the assumption that the discretization-induced transport cost is dominated by local fluctuations: This condition can be naturally satisfied for distributions with sufficiently fast-decaying tails (including Gaussian families), where distant contributions to the quadratic transport cost are reasonably suppressed. In contrast, for distributions exhibiting heavy tails or slow decays, non-local contributions from the tail region may modify the scaling behavior of the $W_{2}$ distance in the limit $d\to 0$, potentially requiring corrections including the higher-order terms in $d$. The present analysis therefore characterizes a universal local correspondence, whose extension to heavy-tailed distributions is left for our future work.
@@
Through the discussions about Gaussian families and its generalization, we now see that the extension of $\delta\xi$ to random vector is not merely a matter of average, but is structurally necessary: The trace structure appearing in the $W_{2}$ of Eq. (24) can only emerge in quadratic form of the KL divergence, if the outer product of $\delta\xi$ is effectively proportional to the identity matrix $I$, i.e., $\delta\xi\delta\xi^{\mathsf{T}}\propto d^{2}I_{f}$. Since the outer product of a single vector is necessarily rank one, while the required structure is full rank with $f$ degrees of freedom, an extension to a probabilistic ensemble of $\delta\xi$ is unavoidable for systems with $f\geq 2$.
@@
We also comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence. While the KL divergence itself is invariant under reparametrizations on the statistical manifold, Eqs. (24)-(25) suggest that its expectation under $\rho$ depends on the trace of Fisher metric $\Omega$. This does not represent a contradiction, but can be reasonably understood from the following two aspects. (i) Taking the expectation $\mathbb{E}_{\rho}$ in Eq. (25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector (the associated probability density $\rho$ typically reflects a modeling choice related to discretization or coarse-graining, depending on the problem under consideration). (ii) In general, when higher-order terms are present, the KL divergence expanded up to second order in $\delta\xi$, yielding a quadratic form, is explicitly governed by the Fisher metric $\Omega$ associated with the chosen parameter $\xi$. In such cases, once combined with the discretization scale and the extrinsic expectation over parameter variations, the resulting numerical value depends on the chosen parameter and Fisher metric. From these viewpoints, the explicit expression $d^{2}\mathrm{Tr}(\Omega)/12$ generally depends on the chosen parameter, yet the correspondence between the 2-Wasserstein transport cost and the expected KL divergence should be understood as an invariant structural relation. The exception is the Gaussian families considered in the UCN analysis, where (i) the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and (ii) the natural parameterization and the standard quadratic transport cost lead to $a=1$. While this value of $a$ is not mathematically unique, it corresponds to a canonical choice in which KL divergence on the statistical manifold and the transport cost induce the same local quadratic form, so that no additional rescaling is required. In more general settings, different choices for parameters typically lead to different numerical prefactors. Nevertheless, the universal correspondence $W_{2}^{2}\propto 2\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs. The explicit prefactor depends only on the extrinsic identification between the discretization scale and infinitesimal parameter variations.
Conceptual Position
In light of the above remarks, it is instructive to clarify the conceptual position of the present approach. The central aim of this work is to clarify how discretization-induced nonlinearity (UCN), inherent to classical discrete systems, can be meaningfully characterized from an information-geometric perspective. At first sight, this objective may appear ill-posed: a continuous probability distribution and its discretized counterpart generally belong to different measure classes and possess incompatible supports, so that their discrepancy cannot be directly quantified within a single statistical manifold using conventional divergences such as the KL divergence. In this sense, a direct information-geometric comparison between continuous and discrete distributions is fundamentally obstructed.
To overcome this difficulty, we do not attempt to directly measure the discrepancy between continuous and discretized distributions themselves. Instead, we relax the original problem and focus on the geometric distortion induced by the discretization procedure viewed as an extrinsic projection from a continuous configuration space onto a discrete lattice. While this projection operation is not definable intrinsically on a statistical manifold, its effect can nevertheless be quantified by introducing an appropriate transport metric on the underlying sample space. The key idea is then to interpret the resulting transport cost by pulling it back to the statistical manifold through infinitesimal parameter variations.
Under discretization, the lattice scale $d$ and infinitesimal parameter variations $\delta\xi$ capture a common notion of scale dependence: On the statistical manifold, $\delta\xi$ controls the distinguishability of nearby distributions, while on the level of the underlying sample space, $d$ governs the transport cost that measures geometric distortions between their supports. To consistently quantify this scale-dependent correspondence on the transport side, its cost function should encode the geometric information controlling the infinitesimal distinguishability on the statistical manifold, i.e., employing a quadratic cost canonically consistent with the local Fisher metric. From this correspondence, $d$ is not merely a numerical resolution parameter, but is explicitly linked to $\delta\xi$.
The resultant 2-Wasserstein distance naturally provides a differential, second-order characterization of the distortion induced by the projection, admits a well-defined limit as $d\to 0$, and allows a direct information-geometric interpretation. The $W_{2}$-KL equivalence should therefore be understood not as an artificial identification of unrelated quantities, but as a consistent pullback of an extrinsic geometric operation into an intrinsic, Fisher-metric-based description on the statistical manifold.
In this sense, the present framework offers a natural and minimal route to interpret discretization-induced nonlinearity in terms of information geometry. Possible alternative metrics or higher-order corrections may encode additional geometric structures; nevertheless, under the requirements of differentiability, quadratic approximability, and a controlled discretization limit, the present transport-information-geometric construction emerges as one of the most natural choices.